OpenStudy (anonymous):

A cylinder whose radius is 4 in and whose height is 8 in has its radius doubled. how many times greater is the volume of the larger cylinder than the smaller cylinder.

OpenStudy (anonymous):

@tomhue

OpenStudy (anonymous):

sorry for lateness Cylinder volume formula is: \[V={\pi}r^2h\] It's the area of a circle, multiplied by the height. Your values for the small cylinder (we'll label it "a") are r=4 and h=8. Va=pi*4^2*8 Va=pi*16*8 Va=128pi Now, we want to double the radius to get the larger cylinder (we'll label it "b"). r=8 and h=8. Vb=pi*8^2*8 Vb=pi*64*8 Vb=512pi How many times greater is the volume of the larger cylinder than the smaller cylinder? This is simple division. Vb/Va=your answer.